# MELA: Multilingual Evaluation of Linguistic Acceptability

### v2

@@ Abstract
In this work, we present the largest benchmark to date on linguistic acceptability: Multilingual Evaluation of Linguistic Acceptability—MELA, with 48K samples covering 10 languages from a diverse set of language families. We establish LLM baselines on this benchmark, and investigate cross-lingual transfer in acceptability judgements with XLM-R. In pursuit of multilingual interpretability, we conduct probing experiments with fine-tuned XLM-R to explore the process of syntax capability acquisition. Our results show that GPT-4 performs on par with fine-tuned XLM-R, while open-source instruction-finetuned multilingual models lag behind by a notable gap. Cross-lingual transfer experiments show that unlike semantic tasks, in-language training data is crucial in acceptability judgements. Results of probing indicate that training on MELA improves the performance of XLM-R on sytax-related tasks. Our dataset will be made publicly available upon acceptance.
@@
Ziyin Zhang, Yikang Liu, Weifang Huang, Junyu Mao, Rui Wang, Hai Hu Dept. of Computer Science and Engineering, Shanghai Jiao Tong University School of Foreign Languages, Shanghai Jiao Tong University School of Arabic Studies, Beijing Foreign Studies University {daenerystargaryen;yikangliu;huangweifang;wangrui12;hu.hai}@sjtu.edu.cn maojunyu@bfsu.edu.cn
@@ 1 Introduction
As a core linguistic competence, it has been argued in the literature of Chomskyan generative syntax that much if not all of such syntactic competence is innate Chomsky (1965). That is, human brains are born with such knowledge already wired in. If the “innate” hypothesis were to be true and linguistic competence were unique in humans, it would naturally follow that any language model—with no “innate” linguistic knowledge to begin with—cannot be taught to acquire certain key linguistic competence.
There have been many attempts in computational linguistics and cognitive science to investigate this hypothesis, directly or indirectly, using either a data-driven approach, where examples created by theoretical linguists in published textbooks are collected, e.g., CoLA—Corpus of Linguistic Acceptability (Warstadt et al., 2019), or a theory-driven approach, where minimal pairs targeting specific syntactic phenomena are generated semi-automatically via some template (Warstadt et al., 2020; Xiang et al., 2021; Hu et al., 2020a).
Recently there has been growing interest in expanding the data-driven paradigm into other languages. For instance, CoLA-style datasets have been proposed in Russian (Mikhailov et al., 2022), Italian (Trotta et al., 2021), and Chinese (Hu et al., 2023). However, to date there are no multilingual benchmarks in this area that can be used to systematically test such abilities of multilingual models.
On the other hand, recently introduced evaluation benchmarks for Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023) have mostly focused on application-driven tasks such as world knowledge and commonsense reasoning (Hendrycks et al., 2021; Srivastava et al., 2022), math reasoning (Cobbe et al., 2021), and code generation (Zhang et al., 2023). Few works, however, have investigated these models from a purely linguistic aspect.
To address these gaps, we introduce MELA—Multilingual Evaluation of Linguistic Acceptability, the first large-scale multilingual acceptability benchmark with 48k examples covering 10 languages from a diverse set of language families. Data in four languages are from existing benchmarks mentioned above, and we complement them with newly collected data in six languages. Examples of MELA are demonstrated in Table 1. Following the CoLA tradition, all sentences in MELA are hand-written by linguists in respective languages, taken from textbooks, handbooks and journal articles in theoretical syntax, except for a small fraction of Russian sentences from Mikhailov et al. (2022).
We come up with three possible usages of MELA. In this work, we make a preliminary exploration in the following three directions:
@@
We probe the syntax capacity of MELA-finetuned XLM-Rs on syntax-related probing tasks, which indicates that XLM-R acquires syntax knowledge from the linguistic judgment task.
In the rest of this work, we first review relevant literature in §2, and then describe how we construct our benchmark MELA in §3. Next, we apply MELA as an evaluation benchmark for LLMs in §4. We investigate cross-lingual transfer in §5 and multi-lingual fine-tuning in §6. Finally, we probe the XLM-Rs trained on MELA for their syntax-related capacity in §7.
@@ 2.1 Linguistic Acceptability
As we mentioned in §1, currently there exist four large-scale linguistic acceptability datasets: CoLA (Warstadt et al., 2019), ItaCoLA (Trotta et al., 2021), RuCoLA (Mikhailov et al., 2022), and CoLAC (Hu et al., 2023), all of which are annotated by expert linguists, while CoLAC also comes with an additional set of crowd labels.
Another line of work in linguistic acceptability is based on semi-automatic construction of example sentences, usually in minimal pairs. They compare the probabilities that language models assign to these sentences (Warstadt et al., 2020; Xiang et al., 2021), sometimes focusing on specific syntactic issues such as agreement (Varda and Marelli, 2023). A recent work also collects acceptability data in six Scandinavian languages (Nielsen, 2023), where the unacceptable examples are automatically generated by removing or swapping words in sentences from the Universal Dependency project.
In this work, we follow the CoLA style when building our benchmark, so that the unacceptable sentences are manually created by linguists to reflect certain syntactic constraints of the language in question. Compared with automatic methods, a wider coverage of syntactic phenomena is achieved in this way.
@@ 2.2 Multilingual Evaluation Benchmarks
XTREME (Hu et al., 2020b) and XGLUE (Liang et al., 2020) are two of the most popular multilingual evaluation benchmarks. Of the tasks therein, many are constructed by translating English samples entirely or partially into other languages, such as XNLI (Conneau et al., 2018), PAWS-X (Yang et al., 2019), and MLQA (Lewis et al., 2020).
Apart from these NLU benchmarks, the literature has also witnessed an abundance of multilingual generation benchmarks, ranging from summarization (Scialom et al., 2020; Ladhak et al., 2020) to translation (Fan et al., 2021; Goyal et al., 2022). After multitask instruction finetuning was found to unlock cross-task generalization ability in language models (Wei et al., 2022; Sanh et al., 2022), multilingual instruction datasets have also been proposed, represented by Supernatural Instruction (Wang et al., 2022) and xP3 (Muennighoff et al., 2023).
English Chinese Russian Italian German French Spanish Japanese Arabic Icelandic ISO code en zh ru it de fr es ja ar is Train 8551 6072 7869 7801 500 500 500 500 500 500 Dev 527 492 1483 946 402 521 321 693 313 1194 Test 516 931 2341 975 402 521 322 694 313 1194 acceptable% 70.3 66.4 73.2 84.5 75.5 85.0 73.1 80.0 74.7 75.1 len (char) 40.8 11.7 56.2 36.0 49.0 29.0 31.7 16.1 22.2 32.9 len (byte) 40.8 35.0 102.8 36.3 49.6 29.7 32.9 47.7 40.8 36.7 len (token) 10.5 9.5 15.2 9.7 11.5 8.1 8.7 11.1 7.9 9.7
@@ 3 MELA: Multilingual Evaluation of Linguistic Acceptability
MELA consists of more than 48 thousand acceptability samples across 10 languages from a diversity of language families and groups. Specifically, it contains three Germanic languages: English, German and Icelandic, three Romance languages: Spanish, French and Italian, one Slavic language: Russian, one Sino-Tibetan language: Chinese, one Japonic language: Japanese, and one Semitic language: Arabic. Table 1 shows example sentences and properties of each language in MELA. For dataset statistics, see Table 2.
@@ 3.1 Data collection Procedure
We use four existing datasets for four languages in MELA: CoLA (Warstadt et al., 2019) for English, ItaCoLA (Trotta et al., 2021) for Italian, RuCoLA (Mikhailov et al., 2022) for Russian, and CoLAC for Chinese (Hu et al., 2023), each having more than 6,000 data points. Since the out-of-domain samples of RuCoLA are produced by generative models, we additionally collected 1037 Russian samples from The Syntax of Russian (Bailyn, 2011) (with the procedure described below) and add them 50-50 to the development and test sets of the Russian portion to keep a balance between validation-test discrepancy and generalization.
@@
As these books/textbooks and handbook are overviews of the syntax of each language, we believe they cover a wide range of linguistic phenomena in these languages, and can therefore serve as a good resource to evaluate language models’ overall ability to distinguish acceptable sentences from unacceptable ones.
@@ 3.2 Resulting Corpus and Data Split
The resulting corpus contains more than 48k example sentences in 10 languages.
For Italian and Chinese, we use the original train/dev/test splits of ItaCoLA and CoLAC, and for CoLAC we use the crowd label following Hu et al. (2023) (see Appendix C for the alternative). For English and Russian, we keep the training splits of CoLA v.1.1 and RuCoLA, and use their in-domain development sets as our validation sets, and their out-of-domain development sets as our test sets.
For the six low-resource languages, we randomly sample 500 sentences from each of these languages to construct a training set, and divide the remaining sentences equally between validation and test sets.111We experimented with another split of these data and observe similar results in all experiments that follow.
@@ 3.3 Evaluation Metric
Following previous works in linguistic acceptability, we evaluate the performance on MELA by Matthews Correlation Coefficient (MCC, Matthews, 1975), which is a measurement of similarity between binary distributions taking values from -1 to 1 and always yielding 0 for any two uncorrelated distributions, regardless of class imbalance.
@@ 3.4 Comparison with Other Multilingual Benchmarks
We note that all samples in MELA are constructed individually in each language. While some early multilingual benchmarks opt to translate English sentences into other languages to obtain parallel samples (Conneau et al., 2018; Lewis et al., 2020), this approach does not suit our case. Firstly, as Clark et al. (2020) argue, translation introduces artifacts into multilingual benchmarks and often results in translationese. Secondly, the task of linguistic acceptability is highly language-dependent, and syntactic phenomena in one language most likely cannot be captured in another language through translation.
model size examples en zh it ru de fr es ja ar is avg Supervised XLM-R 550M - 70.65 55.20 53.97 50.04 37.60 22.46 45.89 44.90 30.59 35.39 44.67 Open-sourced, instruction-finetuned BLOOMZ 7.1B - -2.28 9.99 -1.34 -1.60 -0.90 -3.20 -0.91 -1.86 7.50 3.46 0.88 BLOOMZ 7.1B in-lang. 7.74 17.63 4.87 -0.25 -0.14 -0.34 7.05 3.81 -1.85 -1.92 3.66 BLOOMZ 7.1B en 7.74 12.53 2.66 -0.47 3.86 -1.52 6.30 2.31 -2.36 -3.92 2.71 mT0 13B - 7.32 20.13 10.83 1.95 10.28 0.39 9.32 13.71 0.04 4.95 7.89 mTk 13B in-lang. 39.13 32.18 18.26 11.83 9.91 13.09 24.42 22.45 12.72 15.54 19.95 mTk 13B en 39.13 31.48 12.12 14.92 16.46 12.81 15.77 15.17 6.34 11.21 17.54 Baichuan2-Chat 13B - 13.46 15.78 7.07 13.29 5.77 3.34 16.43 13.85 5.76 -0.98 9.38 Baichuan2-Chat 13B in-lang. 27.26 25.89 13.14 7.23 6.78 6.68 16.43 17.87 3.04 0.94 12.52 Baichuan2-Chat 13B en 27.26 14.88 7.44 1.97 2.76 9.54 13.77 10.19 3.04 -1.05 8.98 Close-sourced GPT-3.5 - - 37.16 30.34 29.43 17.88 29.51 25.59 49.23 31.71 10.24 5.97 26.71 GPT-3.5 - in-lang. 67.00 45.64 38.46 24.47 27.29 23.63 59.76 38.71 18.42 14.60 35.80 GPT-3.5 - en 67.00 15.22 13.99 8.11 13.44 13.86 38.60 16.61 5.28 3.76 19.59 GPT-4 - - 69.31 50.75 35.57 37.87 43.03 32.45 51.52 45.87 16.44 9.88 39.27 GPT-4 - in-lang. 72.29 55.57 51.40 38.31 36.54 35.57 56.16 49.36 17.09 22.60 43.49 GPT-4 - en 72.29 45.49 14.57 -0.94 23.48 12.97 43.66 39.01 3.04 6.62 26.02

### v3

@@ Abstract
In this work, we present the largest benchmark to date on linguistic acceptability: Multilingual Evaluation of Linguistic Acceptability—MELA, with 46K samples covering 10 languages from a diverse set of language families. We establish LLM baselines on this benchmark, and investigate cross-lingual transfer in acceptability judgements with XLM-R. In pursuit of multilingual interpretability, we conduct probing experiments with fine-tuned XLM-R to explore the process of syntax capability acquisition. Our results show that GPT-4o exhibits a strong multilingual ability, outperforming fine-tuned XLM-R, while open-source multilingual models lag behind by a noticeable gap. Cross-lingual transfer experiments show that transfer in acceptability judgment is non-trivial: 500 Icelandic fine-tuning examples lead to 23 MCC performance in a completely unrelated language—Chinese. Results of our probing experiments indicate that training on MELA improves the performance of XLM-R on syntax-related tasks.
@@
Ziyin Zhang†∗, Yikang Liu‡∗, Weifang Huang‡, Junyu Mao⋄, Rui Wang†, Hai Hu‡ † Dept. of Computer Science and Engineering, Shanghai Jiao Tong University ‡ School of Foreign Languages, Shanghai Jiao Tong University ⋄ School of Arabic Studies, Beijing Foreign Studies University {daenerystargaryen;yikangliu;huangweifang;wangrui12;hu.hai}@sjtu.edu.cn maojunyu@bfsu.edu.cn
@@ 1 Introduction
As a core linguistic competence, the ability to tell well-formed sentences from ill-formed ones is one of the first that a good language model should have.
Many corpora and benchmarks have been built to evaluate language models’ syntactic ability, using either a data-driven approach, where examples created by theoretical linguists in published textbooks are collected, e.g., CoLA—Corpus of Linguistic Acceptability (Warstadt et al., 2019), or a theory-driven approach, where minimal pairs targeting specific syntactic phenomena are generated semi-automatically via some template (Warstadt et al., 2020; Xiang et al., 2021; Hu et al., 2020b).
Recently, there has been growing interest in expanding the data-driven paradigm into other languages. For instance, CoLA-style datasets have been proposed in Russian (Mikhailov et al., 2022), Italian (Trotta et al., 2021), and Chinese (Hu et al., 2023). However, to date there are almost no multilingual benchmarks in this area that can be used to systematically test such abilities of multilingual models.
On the other hand, recently introduced benchmarks for Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023) have mostly focused on application-driven tasks such as world knowledge and commonsense reasoning (Hendrycks et al., 2021a; Srivastava et al., 2022), math reasoning (Cobbe et al., 2021; Hendrycks et al., 2021b), and code generation (Chen et al., 2021; Austin et al., 2021; Zhang et al., 2023). Few studies, however, have investigated these models from a more linguistics-oriented aspect.
To address these gaps, we introduce MELA—Multilingual Evaluation of Linguistic Acceptability, the first large-scale multilingual acceptability benchmark with 46k examples covering 10 languages from a diverse set of language families. Data in four languages are from existing benchmarks mentioned above, and we complement them with newly collected data in six languages. Examples of MELA are demonstrated in Table 1. Following the CoLA tradition, all sentences in MELA are hand-written by linguists in respective languages, taken from textbooks, handbooks and journal articles in theoretical syntax, except for a small fraction of Russian sentences from Mikhailov et al. (2022).
We propose three possible usage cases for MELA, and make preliminary explorations in this paper:
@@
We train XLM-R (Conneau et al., 2020) on different language combinations, finding non-trivial cross-lingual transfer performance even between unrelated language pairs, despite the vast difference in the basic syntax of the 10 languages in MELA.
@@
We probe the syntactic capacity of MELA-finetuned XLM-Rs on syntax-related probing tasks, which indicates that XLM-R acquires some syntactic knowledge from finetuning on the acceptability judgment task.
In the rest of this paper, we first review relevant literature in §2, and then describe how MELA was constructed in §3. Next, we use MELA to benchmark several open-source and close-source LLMs in §4. We investigate cross-lingual transfer and multilingual fine-tuning in §5. Finally, we probe the XLM-Rs trained on MELA for their syntax-related capacity in §6.
@@ 2.1 Linguistic Acceptability
As we mentioned in §1, large-scale linguistic acceptability datasets are currently available for four languages: CoLA for English (Warstadt et al., 2019), ItaCoLA for Italian (Trotta et al., 2021), RuCoLA for Russian (Mikhailov et al., 2022), CoLAC for Chinese (Hu et al., 2023), NoCoLA for Norwegian (Jentoft and Samuel, 2023), and JCoLA (Someya et al., 2024) for Japenese. Sentences from these datasets are taken from academic works by theoretical syntacticians and are therefore annotated by expert linguists.111CoLAC also comes with an additional set of crowd labels; Unacceptable sentences in NoCoLA are sourced from grammatical mistakes made by Norwegian learners.
Another line of work in linguistic acceptability is based on minimal pairs, consisting of two near-identical sentences with minimal differences. Language models are expected to assign a higher probability to the acceptable sentence than the unacceptable one. The minimal pair paradigm is adopted to evaluate specific syntactic issues such as subject-verb agreement (Linzen et al., 2016; Marvin and Linzen, 2018; Varda and Marelli, 2023), reflexive anaphora (Futrell et al., 2019; Hu et al., 2020a), negative polarity licensing (Wilcox et al., 2019; Jumelet and Hupkes, 2018), long-distance dependency (Wilcox et al., 2018; Chowdhury and Zamparelli, 2018), and argument structure (Kann et al., 2019; Tjuatja et al., 2023). Following these works, comprehensive benchmarks of minimal pairs are constructed in English resources (Warstadt et al., 2020; Hu et al., 2020b), and then expanded into other languages (Xiang et al., 2021; Song et al., 2022; Someya and Oseki, 2023; Nielsen, 2023).
In this work, we follow CoLA in constructing our benchmark as an initial step towards multilingual evaluation in acceptability judgment, as our goal is to have a wide coverage of syntactic phenomena in the languages selected.
@@ 2.2 Multilingual Evaluation Benchmarks
XTREME (Hu et al., 2020c) and XGLUE (Liang et al., 2020) are two of the most popular multilingual evaluation benchmarks. Of the tasks therein, many are constructed by translating English samples entirely or partially into other languages, such as XNLI (Conneau et al., 2018), PAWS-X (Yang et al., 2019), and MLQA (Lewis et al., 2020).
Apart from these NLU benchmarks, the literature has also witnessed an abundance of multilingual generation benchmarks, ranging from summarization (Scialom et al., 2020; Ladhak et al., 2020) to translation (Fan et al., 2021; Goyal et al., 2022). After the popularization of multitask instruction finetuning in language models (Wei et al., 2022; Sanh et al., 2022), multilingual instruction datasets have also been proposed, represented by Supernatural Instruction (Wang et al., 2022) and xP3 (Muennighoff et al., 2023). We refer to Qin et al. (2024) for a more comprehensive review of recent multilingual resources.
English Chinese Italian Russian German French Spanish Japanese Arabic Icelandic ISO code en zh it ru de fr es ja ar is Trainv1.0 8551 6072 7801 7869 500 500 500 500 500 500 Devv1.0 527 492 946 1405 272 466 295 580 258 899 Testv1.0 516 931 975 2227 273 467 293 581 259 899 Trainv1.1 8551 6072 7801 7869 - - - - - - Devv1.1 527 492 946 1405 100 100 100 100 100 100 Testv1.1 516 931 975 2227 945 1333 988 1561 917 2198 acceptablelen (char) 33.1 10.7 30.0 47.9 39.9 22.9 26.2 14.7 18.1 26.4 len (byte) 34.1 34.3 31.3 95.7 41.5 24.6 28.4 47.0 38.3 31.1 len (token) 10.5 9.5 9.7 15.3 11.4 8.1 8.7 10.9 8.2 9.4
@@ 3 MELA: Multilingual Evaluation of Linguistic Acceptability
MELA consists of more than 46 thousand acceptability samples across 10 languages from a diversity of language families and groups. Specifically, it contains three Germanic languages: English, German and Icelandic, three Romance languages: Spanish, French and Italian, one Slavic language: Russian, one Sino-Tibetan language: Chinese, one Japonic language: Japanese, and one Semitic language: Arabic. Table 1 shows example sentences and properties of each language in MELA. For dataset statistics, see Table 2.
@@ 3.1 Data collection Procedure
We use four existing datasets for four languages in MELA: CoLA (Warstadt et al., 2019) for English, ItaCoLA (Trotta et al., 2021) for Italian, RuCoLA (Mikhailov et al., 2022) for Russian, and CoLAC for Chinese (Hu et al., 2023), each having more than 6,000 data points.222NoCoLA and JCoLA are not included for they are concurrent with this work. Since the out-of-domain samples of RuCoLA are produced by generative models, we additionally collected 1037 Russian samples from The Syntax of Russian (Bailyn, 2011) (with the procedure described below) and add them 50-50 to the development and test sets of the Russian portion to keep a balance between validation-test discrepancy and generalization.
@@
As these books/textbooks and handbooks are overviews of the syntax of each language, we believe they cover a wide range of linguistic phenomena in these languages, and can therefore serve as a good resource to evaluate language models’ overall ability to distinguish acceptable sentences from unacceptable ones.
@@ 3.2 Resulting Corpus and Data Split
The resulting corpus contains more than 46k example sentences in 10 languages.
For Italian and Chinese, we use the original train/dev/test splits of ItaCoLA and CoLAC, and for CoLAC we use the crowd label following Hu et al. (2023) For English and Russian, we keep the training splits of CoLA v.1.1 and RuCoLA, and use their in-domain development sets as our validation sets, and their out-of-domain development sets as our test sets.
For the six low-resource languages, we decide to adopt two splits for two purposes: fine-tuning smaller models such as XLM-R (v1.0) and benchmarking LLMs (v1.1).333Performance of LLMs on two splits of these data are similar (see Table 8). For v1.0, with the purpose for fine-tuning, we randomly sample 500 sentences from each of these languages to construct a training set, and divide the remaining sentences roughly equally between validation and test sets. For v1.1, we reserve 100 samples from each language as the validation set, and keep all the rest of the examples in the test set, thus producing a larger test set which we believe will make the evaluation more stable. See Table 2 for details of the two splits.
@@ 3.3 Evaluation Metric
Following previous works in linguistic acceptability, we evaluate the performance on MELA by Matthews Correlation Coefficient (MCC, Matthews, 1975), which is a measure of similarity between binary distributions taking values from -1 to 1 and always yielding 0 for any two uncorrelated distributions, regardless of class imbalance.
@@ 3.4 Comparison with Other Multilingual Benchmarks
We note that all samples in MELA are constructed individually in each language. While some early multilingual benchmarks opt to translate English sentences into other languages to obtain parallel samples (Conneau et al., 2018; Lewis et al., 2020), this approach does not suit our case. First, the task of linguistic acceptability is highly language-dependent, and syntactic structures acceptable in one language may not be acceptable in another, and thus there is no easy way of translating existing corpora into other languages while keeping the target syntactic phenomena. Second, as Clark et al. (2020) and Hu and Kübler (2021) argue, translation introduces artifacts into multilingual benchmarks and often results in translationese.
model size examples en zh it ru de fr es ja ar is avg Supervised XLM-R 550M - 60.64 54.94 53.53 49.37 26.72 19.04 34.08 29.32 14.12 35.41 37.72 Open-sourced BLOOMZ0 7.1B - 4.49 14.13 4.83 4.77 1.63 7.08 10.12 3.27 8.12 0.00 5.85 BLOOMZ2 7.1B in-lang. -1.11 7.65 5.67 5.38 3.90 5.19 6.76 3.83 6.22 -0.35 4.31 BLOOMZ2 7.1B en -1.11 7.90 4.22 4.74 0.96 4.72 8.07 2.45 4.65 -0.09 3.65 mT00 13B - -5.42 9.68 12.68 5.57 11.33 8.24 2.88 13.20 6.77 1.22 6.62 mT02 13B in-lang. 3.54 8.76 10.93 9.04 5.35 6.66 4.72 12.41 8.95 6.61 7.70 mT02 13B en 3.54 8.06 10.22 10.68 7.17 7.40 6.09 10.46 2.91 4.86 7.14 mTk0 13B - 5.25 -1.49 -3.62 5.90 1.25 3.81 5.82 1.04 1.85 2.59 2.24 mTk2 13B in-lang. 22.74 8.47 10.24 16.66 11.96 9.28 13.34 12.00 4.87 10.93 12.05 mTk2 13B en 22.74 8.36 8.98 15.69 14.54 12.30 9.28 10.92 6.52 5.99 11.53 Baichuan2-Base0 13B - 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 Baichuan2-Base2 13B in-lang. 46.11 47.36 24.01 28.84 13.40 17.41 21.95 20.68 13.90 1.81 23.55 Baichuan2-Base2 13B en 46.11 35.16 17.84 25.88 6.42 15.95 16.57 13.48 11.41 -3.80 18.50 Baichuan2-Chat0 13B - 37.15 33.56 10.08 7.93 -6.49 8.41 18.32 11.15 0.00 2.01 12.21 Baichuan2-Chat2 13B in-lang. 41.12 29.25 18.10 19.46 6.46 18.57 20.81 14.18 13.97 -1.51 18.04 Baichuan2-Chat2 13B en 41.12 27.02 12.22 14.11 2.80 9.49 14.62 11.40 7.00 -5.03 13.47 Close-sourced GPT-3.50 - - 64.60 17.01 14.33 18.05 23.01 31.66 24.35 16.61 9.57 4.69 22.39 GPT-3.52 - in-lang. 64.11 25.32 38.66 21.59 21.62 29.52 44.20 21.48 6.19 9.70 28.24 GPT-3.52 - en 64.11 30.25 25.27 24.91 24.54 29.88 37.75 21.70 6.43 0.56 26.54 GPT-4o0 - - 69.05 62.38 53.01 55.24 36.61 37.01 58.13 50.32 29.86 40.63 49.22 GPT-4o2 - in-lang. 72.14 59.01 54.86 59.17 39.66 37.19 61.36 52.03 32.38 43.64 51.14 GPT-4o2 - en 72.14 54.77 52.51 52.96 39.20 39.50 52.61 47.98 30.08 31.61 47.34